If I have an equilateral triangle with a hypotenuse of 82 cm, what is the length of the two equal sides?
Answer by dolly(163)

Let the 2 equal sides measure x cm
Then by Pythagorean theorem,
x^2 + x^2 = 82 ^2
==> 2x^2 = 6724
==> x^2 = 6724/2
= 3362
==> x = sqrt (3362)
= 57.98 cm
So each equal side measures 57.98 cm